OpenStudy (anonymous):

If you do 1,000 J of work on a system and remove 500 J of heat from it, what is the change in its internal energy

OpenStudy (ybarrap):

Doing 1000 J of work means that the system is doing negative work. Removing 500 J of heat from the system means that the system is losing heat: U=Q-W=-500-(-1000)=-500+1000=500J The internal energy (U) increases by 500 J, since the quantity U is positive.

OpenStudy (anonymous):

thank you I had a feeling it might be 500 J just wasn't sure :)

OpenStudy (ybarrap):

One important correction (which doesn't change our answer). If heat is added to the system q > 0. Similarly, if work is done to the system, w > 0. Both of which increase internal energy. So \(U=Q+W\) NOT \(U=Q-W\). However, because I indicated that work done by the system was negative, we still got the right answer, but I think its much easier to think of work \(\color{red}{added}\) and heat \(\color{red}{added}\) \(\color{red}{to}\) the system as being positive: $$ U=Q\color{red}{+}W\\ =-500+1000=500~J $$ Now (and I hope that this does not to make things too confusing) If we know we have a reversible system, then \(w=-\int PdV\). Which makes internal energy equal to $$ U=Q-\int_{v_1}^{v_2} PdV $$ (which is where the negative sign comes in) Where if work were done to the system then \(v_2 < v_1\), which would increase internal energy, because then the integral part of U would be positive (i.e. \(-\int_{v_1}^{v_2} PdV > 0\))